Operations Report – December 10th

Crew 184 Operations Report 10/12/2017

SOL: 9

Name of person filing report: HuntJ

Non-nominal systems: Generator non-functional, repaired and returned to operations in afternoon. (Operational Again!!!)

Notes on non-nominal systems: Mission Support found a solution with the generator today. It turns out that the ground and positive connectors to the battery were loose. Tightening them and a voltimeter check proved the battery was just fine. One more attempt at using jumper cables and the generator is back and running! Forward plan now is to leave it on until tomorrow morning (with fingers crossed that it stays on through the night). A decision will be made tomorrow morning on whether to turn the generator back off or leave it running.

Generator (hours run): Back on around 2:00pm. Running 3 hours and still going!

Solar— SOC @ 4:00pm: 99%

Diesel – 100%

Propane – 70 Percent Volume

Ethanol Free Gasoline (5 Gallon containers for ATV) – 1.5 Gallons

Water (trailer) – 600. gallons

Water (static) – 250 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 47 gallons

Toilet tank emptied: No

ATV’s Used: None,

Oil Added: No

ATV Fuel Used: None

# Hours the ATVs were Used today: 00:00 Hours

Notes on ATVs: No ATVs used today but were turned on for status checks. All nominal except for the Yamaha 300 which was determined a few days ago to have a dead battery.

Deimos rover used: Yes

Hours: 01:00 Hours

Beginning charge: 100%

Ending charge: 88.8%

Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Yes

Hours: 01:00

Beginning charge: 100%

Ending charge: 51%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 01:00 Hours

Beginning charge: 100%

Ending charge: 23%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 00:30 Hours

Beginning charge: 100%

Ending charge: 59%

Currently charging: Yes

HabCar used and why, where? NO

General notes and comments: Power is back and we’re back in sim!!

Summary of internet: 1.0 GB of Bandwidth

Summary of suits and radios: Radios are nominal after getting charged again.

EVA Suits:

EVA 6: Fan does not turn on.

EVA 8: Latches that secure the back panel are loose. Engaging them again does not change this. Concern is that back panel would fall off during EVA.

We will try to troubleshoot some of these issues and will update with changes.

Summary of Hab operations: Nominal

Summary of GreenHab operations: See Greenhab report

Summary of ScienceDome operations: Currently being cleaned

Summary of RAM operations: Not Operational

Summary of health and safety issues: All Healthy

Questions, concerns and requests to Mission Support:

Request more ATV Fuel. Can probably get through 2 more EVAs before exhausting remaining reserves.
Non burnable trash in airlock ready for pickup.

Operations Report – December 7th

SOL: 6

Name of person filing report:  Thomas Horn

Non-nominal systems:

1.    Generator – see previous ops reports for details

2.    Batteries – After turning off the generator batteries ran from 100% to 25% overnight.  Per Mission Control this is off nominal and implies the batteries have become degraded.  We will continue to monitor throughout the day

Generator—

1.    Turned on at 9am.

Solar—

29% @ 7am
13% @ 9am
30% @ 10am

37% @ 11am
35% @ 11:40am

35% @ 4pm

100% @ 6pm

Diesel – 45% full

Propane –  71%

Gasoline (2x 5 Gallon containers for ATV) –  2.0 gallons remains

Water (trailer) –  600 gallons

Water (static) –  60% remaining

Trailer to Static Pump used –  No

Water (loft) – Static to Loft Pump used –  Yes

Water Meter:  48 gallons

Toilet tank emptied:  Yes

ATV’s Used:  2

Oil Checked:  No

Atv # Fuel Used Gals:  0.5 gallons approx.

Tires Status:  Yamaha ATV #3 Back Right Wobbly

# Hours the ATVs were Used today: 1h

Rovers (changed from Deimos to give Rover general status) used:  Yes – Spirit

Hours: 1h

Beginning charge:  100%

Ending charge: 100%

Currently charging: Yes

HabCar used?  No

Notes and Comments:

Summary of  Hab operations:

1.    CAT construction machine removed by Mission Control

2.    Burnable Trash removed.  Non-burnable remains in Airlock

Summary of GreenHab operations:  Nominal – see greenhab report for details

Summary of Science Dome operations:  None

Summary of health and safety issues: None

Questions, concerns and requests to Mission Support:

1.     Deliver additional ATV fuel

2.     Return Crew 184 Flash Drive

3.     Remove Non-Burnable trash from Rear Airlock

4.     News Crew absent today – confirm if we will have any external media visits during our mission

Operations Report – December 6th

Crew 184 Operations Report  06/12/2017

 

SOL: 5

 

Name of person filing report: Josh Hunt

Non-nominal systems: Generator: Generator went off again overnight. Upon inspection of the generator, the fault light on the generator display was red-orange. Mission support was called and they were able to start up the generator again. Source of anomaly is believed to be the fuel itself which is still running on the “summer blend” and needs the “winter blend”. Winter blend has been ordered by Mission Support. Will log the time when the winter blend fuel arrives. Short-term forward plan is to turn off the generator at 10:00 pm tonight and turn beck on before 8:00 am tomorrow. 

                                   Hab Leak: EVA crew identified a water leak at the bottom of the Hab below the airlock porch. Mission support identified that this is a leak from the Hab Kitchen and is expected. Will monitor in case leak increases in severity in the future.

 

Generator— Hours run: Unknown (See Non Nominal Systems for details)

 

Solar— Morning SOC Charge was 47% 

             Noon SOC Charge was 100% 

             Evening SOC Charge (6:00pm) was 100%

 

Diesel – 46% Full

 

Propane –  72 Percent Volume

 

Gasoline (5 Gallon containers for ATV) –  2.5 Gallons.

 

Water (trailer) –  600 gallons

 

Water (static) –  400 gallons

 

Trailer to Static Pump used –  No

 

Water (loft) – Static to Loft Pump used –  Yes

 

Water Meter: 25 gallons

 

Toilet tank emptied: No

 

ATV’s Used: Yamahas 1 and 2

 

Oil Checked: Yes. All nominal.

 

Atv # Fuel Used Gals: Negligible

 

Tires Status: Yamaha 3 rear right tire is still wobbly. ATV is moved away for the time being and does not affect ops. Mission Control will address at their convenience.

 

# Hours the ATVs were Used today: 00:30 minutes

 

Deimos used: Yes

 

Hours: 00:30 minutes

 

Beginning charge: 87.4 %

 

Ending charge: 87.4 %

 

Currently charging: Yes.

 

 

 

HabCar used? No

 

Notes and Comments:

 

Summary of  Hab operations: Generator continues to provide trouble. This is the second night in a row that it has shut down ungracefully. Forward plan tonight is to check SOC at 10:00pm and then shut off the generator. 

 

 

Summary of GreenHab operations: All nominal

 

Summary of Science Dome operations: All nominal other than generator (see Summary of Hab Operations) 

 

Summary of health and safety issues: All nominal

 

 

Questions, concerns and requests to Mission Support: 

  1. Requesting extra fuel for ATVs. This is the same request previously submitted. We have 2.5 gallons of reserve fuel.
  2. Trash pickup from the rear airlock

Operations Report – December 5th

Crew 184 Operations Report 12/05/2017

SOL: SOL 4

Name of person filing report: Akash Trivedi

Non-nominal systems: Water (pump/lines – static to loft), generator
failure, sudden loss of data

Notes on non-nominal systems:  Water: tanks from static to loft not
transferring due to speculated frozen pipe or blockage. Waited until
10am as per protocol to see if frozen pipes thawed. At 10am, still
frozen. Waited until 12pm for sunlight to be in contact with pipes to
see if thaws. After the meeting at lunch, we learned that a plug by
the static tank needed to be seated. This allows heating along the
pipes. Once plugged we waited some time, and pump was operational
again at around 3pm. Anomaly resolved.

Generator: Noticed generator off at 0900 having failed during the
night. Turned on and seemed to be operating nominally. Returned to
check up on it at 1000, and noticed a coolant leak from the radiator.
Generator turned off and leak corrected by tightening the antifreeze
cap. Generator turned back on and checked status on power console –
seems to not be doing the charging as it’s on “absorb” and SOC not
increasing despite minimal hab power usage. After Lunch meeting, we
monitored thegenerator closely and at sunset the SOC returned to 100%.
Will continue to monitor closely, but power is good at this time.

Loss of internet data: Data check was conducted at 0845 when it was
approximately 800MB, and an hour later, this dropped to zero. Reason
unknown.

Generator (hours run):  Unknown due to night time failure

Solar— SOC % (Before generator is run at night) 4% at 0900, 14% at
1000 (when coolant leak discovered), static 9% on “absorb charging”
with minimal hab usage and generator powered on

Diesel – 48%

Propane –  79 percent volume

Ethanol Free Gasoline (5 Gallon containers for ATV) – 1.25 gallons

Water (trailer) – 600 gallons

Water (static) – 450 gallons

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 46 gallons

Toilet tank emptied: Yes

ATV’s Used: Honda ATV, Yamaha 1

Oil Added? No

ATV Fuel Used: 1.25 gallons

# Hours the ATVs were Used today: 00:45

Notes on ATVs: Yamaha 3 Rear right tire is still wobbly, EVA team
assessed and determined it will require. Photos will be provided in
overall Sol 4 package for awareness and review.

Deimos rover used: Used

Hours: 00:45

Beginning charge: 100%

Ending charge: 100%

Currently charging:Yes

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:  Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments: Multiple non-nominal systems!!!

Summary of internet: Sudden anomalous data loss – see above

Summary of suits and radios: All nominal based on last usage

Summary of Hab operations: Loft water tank almost empty, multiple
non-nominal systems

Summary of GreenHab operations:  All nominal

Summary of ScienceDome operations: All nominal

Summary of RAMM operations: Not Operational

Summary of health and safety issues: No problems now that water, heat,
and power are back.

Questions, concerns and requests to Mission Support: See above on
non-nominal systems. Lack of training and relevant manuals for
troubleshooting significant issues in mission critical systems. Poor
insulation of static tank to loft tank pipes – could lead to pipe
cracking

Operations Report – December 4th

SOL: SOL 3

Name of person filing report: HuntJ

Non-nominal systems:

  1. Wi-Fi Modem (lack of range).  The wifi in a roughly 5 feet circle from the router, any farther and signal drops off line.  We have tested this and created a signal service map by carrying laptops around the habitat to determine when it picks up signal.  No known fix, will operate internet within this range for now.
  2. HAB Eastern Wall has sustained wind damage, panels are beginning to peel off wall. When wind blows the panels swing and hit support beam causing HAB to vibrate.  In order to fix we need HAB material adhesive delivered, after which we will go EVA and patch the damage,
  3. ATV Yamaha 250 #3 Back Right Wheel is loose, observed to be wobbling during driving.  We will assess a fix during tomorrows EVA – it is possible it simply needs to be tightened, if there is actual mechanical damage then we will move the ATV and declare it off limits.
  4. 3 broken old radios.  All newly delivered radios are functional (thank you supply drop).  Assessing broken radios to determine a fix, if none identified will mark as trash and place in a ‘broken equipment bin’.  4 EVA earpieces were declared nonfunctional, though this may have been due to incorrect settings, will continue to investigate during tomorrows EVA

Generator (hours run): 15.4 Hours

Solar— SOC % : @ 7:34 AM: 65%

                            @ 8:00 AM: 100 % (Generator stopped at this time)

@ 5:34 AM: 42% (Generator started at this time)

Diesel –  48% Full.

Propane –  79 Percent Volume .

Ethanol Free Gasoline (5 Gallon containers for ATV) –   gallons. 2.5 Gallons. One 5 gallon tank empty, second tank half full. Request for extra fuel.

Water (trailer) –  600 Gallons

Water (static) –   500 gallons (We have water again!!)

Trailer to Static Pump used –  Yes

Water (loft) – Static to Loft Pump used –  Yes

Water Meter:  35 gallons

Toilet tank emptied: No

ATV’s Used: 3 Yamaha 250’s and 1 Honda ATV used.

Oil Added? No

ATV Fuel Used:  7.5 gallons used to fill all ATVs to full, though the ATVs did not start with full tanks for our mission.

# Hours the ATVs were Used today: 02:30 hours

Notes on ATVs: See above for wobbly ATV wheel issue.

Deimos rover used: Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

 

Spirit rover used:  Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

 

Opportunity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments: Trailer to Static water tank was used today. Pump was dirty and needed cleaning and we did so. Actual process was nominal. We are now full on water again. When inspecting trailer tank water the front tank had a layer of FOD and detritus at the water line. We will keep watch of filter usage in case this increases our burn through rate on filters.

Summary of internet:  Wifi was fixed this morning, though range is limited (see above). Please send hardware manual for the router and modem to allow us to diagnose issues and repair anything that comes up in the future (for example – to allow us to diagnose what certain light combinations signify)

Summary of suits and radios: See above

Summary of Hab operations: Brick Communication Experiment / CPR experiments successful, exercise on all crew. EVA Outer Door repaired.

Summary of GreenHab operations: No fire extinguisher observed – please deliver in future supply drop.

Summary of ScienceDome operations: Soil Salinity experiments (previous crew), dehydrating dirt for use as counterbalance and weights in exercise equipment (filling bags and medicine balls)

Summary of RAMM operations: Not Operational

Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support:

Need extra fuel for ATVs. Of the two red fuel cans, one is empty, one is half full.

Pre-Mission food inventory is complete, will be sent as it’s own report. Please reference that report for requested supplies to address gaps in current inventory. Please give ETA for supply drop so we can adequately ration food.

Thank you for your assistance!

Operations Report – December 3rd

Crew 184 Operations Report  03/12/2017

SOL: SOL 2

Name of person filing report: HuntJ

Non-nominal systems: Toilet Tank, Water (static)

Notes on non-nominal systems:  Toilet Tank: Tank was particularly filled up during purge and caused all crew to have to move upstairs for duration of purge. Workaround is to purge every other day of ops (e.g Sol 4 is next purge) to make upkeep easier on all.

We also request a resupply of the TB-Cide Quat Cleaner Deordorizer/Disinfectant. We are out after the last purge.

 

                                                    Water (Static): Water is at 50 Gallons which is alarmingly low. Our trailer water is at 1000 gallons but we need the means to transfer the water. Request to get Trailer water moved to Static water or permission and high level instructions to do it ourselves.

Generator (hours run): 13.2 Hours

Solar— SOC % (Before generator is run at night) Not nightfall as of submission, but will report this info in the next report.

Diesel –  50% full

Propane –  Did not check (No EVA), assume to be 79 percent volume due to past readings and trends.

Ethanol Free Gasoline (5 Gallon containers for ATV) –   gallons. 5 gallons

Water (trailer) –   1000 Gallons

Water (static) –   50 Gallons

Trailer to Static Pump used –  No

Water (loft) – Static to Loft Pump used –  Yes

Water Meter:  20 gallons (46 once Pump used)

Toilet tank emptied: Yes

ATV’s Used: None

Oil Added? No

ATV Fuel Used:  None

# Hours the ATVs were Used today: 00

Notes on ATVs: N/A

Deimos rover used: Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:  Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments:

Summary of internet:  Still good so far.

Summary of suits and radios: All nominal

Summary of Hab operations: Static Water Tank almost empty, need means to transfer Trailer water over to Static.

Summary of GreenHab operations:  All nominal

Summary of ScienceDome operations: All nominal

Summary of RAMM operations: Not Operational

Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support: Is there a way we can check to see how much internet we are using during the day? We tried the link provided in a different log template but it didn’t take us to anything.

Operations Report – December 2nd

Crew 184 Operations Report 02/12/2017

SOL: SOL 1

Name of person filing report: HuntJ

Non-nominal systems: All Nominal

Notes on non-nominal systems: N/A

Generator (hours run): N/A (Arrived during day when already off)

Solar— SOC % (Before generator is run at night) 80%

Diesel – 50% Full.

Propane – 80 Percent Volume .

Ethanol Free Gasoline (5 Gallon containers for ATV) – gallons. 5 gallons

Water (trailer) – 1000 Gallons!!!.

Water (static) – 500 gallons

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – No

Water Meter: 46 gallons

Toilet tank emptied: No

ATV’s Used: Honda, 300, 350.1, 350.2, 350.3

Oil Added? No

ATV Fuel Used: Negligible (Test drive only)

# Hours the ATVs were Used today: 00:20 hours

Notes on ATVs: They’re running great!!

Deimos rover used: Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where? No

General notes and comments: Happy to see the Red Planet!

Summary of internet: So far so good!

Summary of suits and radios: All nominal

Summary of Hab operations: All nominal

Summary of GreenHab operations: All nominal

Summary of ScienceDome operations: All nominal

Summary of RAMM operations: Not Operational

Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support: Everything is Awesome!!

Operations Report – December 1st

Crew 183 + Operations Report 1 Dec 2017

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: Phone is still not functional, Low water pressure in Hab

Notes on non-nominal systems: HudgesNet was here all day but unable
to repair the phone. They will be working on it again on Monday. The
filter than normally clogs up the water line in the Hab is missing,
but two other filters were cleaned (Bathroom and kitchen sink) and
water pressure is once again nominal throughout the Hab. A new master
filter will be installed to forever end this problem, which was
probably why the old water heater died.

Generator (hours run): off 8 am on 6 pm

Solar— SOC 94% (Before generator is run at night)

Diesel – 50 %

Propane – 100 psi

Ethanol Free Gasoline (5 Gallon containers for ATV) – not known gallons.

Water (trailer) – 0 gallons. Curtis took the water trailer to town to fill it.

Water (static) – not known gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: not known

Toilet tank emptied: no

ATV’s Used: yes

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 2

Notes on ATVs: nothing to report

Deimos rover used: yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: no

Sojourner rover used: ASSIGNED TO DIRECTOR, Operational

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Operational

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: Operational

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: Not operational

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where? No

General notes and comments: Remember that for the first month of use,
the new rovers will not hold a charge as well as Deimos. We took the
three new ones on a 5 mile round trip and they were each reduced to
about 75%. It takes a while for the batteries to perform at their
peak.

I’d like to say that the rovers were made operational by Brandon and I
but it was actually Chris who did a majority of the work and taught me
a lot about the right tools for the job. Who is Chris? Well, he’s
the guy who was here to fix my phone. It took so long he offered to
help.

The static tank was insulated today. A new heater was installed and
the pipe to the Hab was insulated as well. Shannon’s water tank was
delivered and will be installed tomorrow.

Water heater: After determining that the new water heater would not
fit in the spot where the old one was mounted, it was moved to a spot
on the lower deck just to the right of the ductless propane heater.
It was operational, but would not function correctly until the water
pressure was restored. It is currently connected to the heater’s
propane line. I’m assuming Curtis will reconnect the heater as soon
as he has the supplies.

Summary of internet: not known

Summary of suits and radios: nothing to report

Summary of Hab operations: Water system insulated on the Hab’s
exterior. New hot water heater installed in new location. Water
pressure restored. Water trailer to be refilled.

Summary of GreenHab operations: The thinned plants look mostly good.
One of the tomatoes I tried to transplant does not look good. The
rest look fine. Door was open for a few hours today but it was cold
enough that it stayed mostly warm and not hot.

Summary of ScienceDome operations: Brandon concluded his work.
Nothing more to report.

Summary of RAMM operations: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: Nothing, thank you.

Operations Report – November 30th

Crew 183 + Operations Report 30NOV2017

SOL: n/a

Name of person filing report:  Shannon Rupert

Non-nominal systems: Nothing new to report

Notes on non-nominal systems:  n/a

Generator (hours run): off 8 am on 6 pm

Solar— SOC 90% (Before generator is run at night)

Diesel –  50 %

Propane – 100 psi

Ethanol Free Gasoline (5 Gallon containers for ATV) – not known gallons.

Water (trailer) – 0 gallons.

Water (static) – 250 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter:  not known

Toilet tank emptied: no

ATV’s Used: Honda

Oil Added? No

ATV Fuel Used:  2 Gals

# Hours the ATVs were Used today: 4

Notes on ATVs: nothing to report

Deimos rover used:  yes

Hours:

Beginning charge: 100

Ending charge: 94

Currently charging: yes

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:  Installation in progress

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Installation in progress

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments: Burn barrel not complete. Turns out you
can’t just poke holes in it with a pickaxe.  We have a plan for
tomorrow.  Also, contractors will be onsite all day tomorrow.

Summary of internet:  not known

Summary of suits and radios: nothing to report

Summary of Hab operations: nothing to report

Summary of GreenHab operations: Door was opened this afternoon.
Thinned the babies.  Used 4 gallons of water for watering and prepping
the soil for herbs.

Summary of ScienceDome operations:  Brandon continues his work.
Nothing more to report.

Summary of RAMM operations: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: Nothing tonight, thank you.

Operations Report – November 29th

Crew 183 + Operations Report 29NOV2017

SOL: n/a

Name of person filing report:  Shannon Rupert

Non-nominal systems: Nothing new to report

Notes on non-nominal systems:  n/a

Generator (hours run): off 8 am on 5:30 pm

Solar— SOC  91% (Before generator is run at night)

Diesel –  50 %

Propane –  100 psi

Ethanol Free Gasoline (5 Gallon containers for ATV) – not known gallons.

Water (trailer) – 0 gallons.

Water (static) – 250 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  not known

Toilet tank emptied: no

ATV’s Used: Honda

Oil Added? No

ATV Fuel Used:  2 Gals

# Hours the ATVs were Used today: 3.5

Notes on ATVs: nothing to report

Deimos rover used:  no

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:  Installation in progress

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Installation in progress

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments:  Propane was delivered today.  A leak was
found in the line of the small tank at the director’s house, and it
was fixed.  We began work on the second rover today but then decided
to go to storage and see it we could find better tools.  We found a
socket set that was the right size!  We will get back on it tomorrow.
We also brought a 55 gallon drum from town.  Tomorrow I will pound
holes in it and set it up as a burn barrel.  I will also need to clean
up the old burn barrel area as non-burnable garbage was burned and it
will need to be picked up and taken for disposal.  Lucky me!

Summary of internet:  not known

Summary of suits and radios: nothing to report

Summary of Hab operations: Three new smoke detectors arrived today.

Summary of GreenHab operations: Door was opened this afternoon.  Two
gallons of water was used in the new planters.  I did not water the
plants today.

Summary of ScienceDome operations:  Brandon continues his work.
Nothing more to report.

Summary of RAMM operations: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: Thanks for all of
the help with the rover installation question last night.  No
questions tonight.

Copyright © The Mars Society. All rights reserved. | Main Site